OpenStudy (anonymous):

HELP ME PLEASE>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>Point X is located at (2, −6) and point Z is located at (0, 5). Find the y value for the point Y that is located one over five the distance from point X to point Z. −3.6 −3.8 −4 −4.2

OpenStudy (anonymous):

Please someone help me

OpenStudy (anonymous):

@navk any idea?

OpenStudy (anonymous):

Have you tried using the section formula for this question?

OpenStudy (anonymous):

No i havent?

OpenStudy (anonymous):

Okay this question can be solved by it: Point Y(x, y) divides the line XZ in the ratio 1:5. So by section formula, y = (1 * 5 + 5 * -6)/(1 + 5)\[y =\frac{ 1\times5+5\times-6}{1+5}\] That gives you y = -25/6 = -4.2

OpenStudy (anonymous):

@navk YOU ARE AMAZING................ Any way you could help me with one more??? please?

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

@navk Point P is located at (−2, 7) and point R is located at (1, 0). Find the y value for the point Q that is located two over three the distance from point P to point R. 4.9 4.7 2.5 2.3

OpenStudy (anonymous):

Here point Q divides PR in the ratio 2:3 so m = 2 and n = 3. Section formula is\[y = \frac{m \times y_2 + n \times y_1}{m + n}\] where y2 and y1 are the y-coordinates of point P and R. Can you solve it from here on your own?

OpenStudy (anonymous):

@navk I honestly cannot... i need to work on this i know /;

OpenStudy (anonymous):

We got m = 2, n = 3 now y2 is the y coordinate of point R so y2 = 0 and y1 is coordinate of point P so y1 = 7. Plugging these values in the section formula we get y = (2*0 + 3*7)/(2 + 3) = 21/5 = 4.2 The answer should be 4.2 but its not given in the choices? Can you check the choices?

OpenStudy (anonymous):

4.9 4.7 2.5 2.3

OpenStudy (anonymous):

@navk thats the choices...... im thinking 4.7?

OpenStudy (anonymous):

I'm pretty sure it is 4.2 maybe the 4.7 was meant to be 4.2

OpenStudy (anonymous):

@navk Thinking i should put 4.7??

OpenStudy (anonymous):

@livingitup5611 let me recheck it once

OpenStudy (anonymous):

@navk Sounds good

OpenStudy (anonymous):

Okay try to put in 4.7 maybe it will work..

OpenStudy (anonymous):

@navk Thank you for all of your help!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

You're welcome :)